0211avana111:

d. What is the power of a lightbulb that has a current of 0.480 A at 120.0 V? (3 points) e. What is the voltage drop across a 40.0 Ω resistor that has a 0.350 A current?

JamesTDG:

For d, you need to find R in the equation R=V/I, since we know that V= 120 and I= 0.48, the equation is R = 120/0.48